OpenStudy (anonymous):

Find y' and simplify y=3x^-2+2x^3+5

OpenStudy (shamim):

U know y'=dy/dx Right?

OpenStudy (shamim):

Differentiation of y with respect to x

OpenStudy (shamim):

Do u know dx^4/dx=4*x^3

OpenStudy (shamim):

dx^-3/dx=-3*x-4

OpenStudy (shamim):

One more thing is d(4)/dx=0

OpenStudy (shamim):

Because differentiation of any constant is equal 0

OpenStudy (unklerhaukus):

\[y=3x^{-2}+2x^3+5\] \[y'=(3x^{-2})'+(2x^3)'+(5)'\] use \((ax^n)'=anx^{n-1}\)